# A microscopic theory of Anderson localization of electrons in random lattices

### v1

@@ II.3 Ward identities, diffusion pole, and generalized Einstein relation
We must know the dynamical electron-hole Green function $\mathcal{G}^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ to determine the electron-hole correlation function and transport properties of the model. It depends on the one-particle Green functions $G^{R}_{\mathbf{k}}(E_{+})$ and. $G^{A}_{\mathbf{k}}(E_{-})$. The solution from the two-particle functions cannot be disconnected from the solution of the one-particle Green function. The one-particle and two-particle functions are connected via Ward identities. There are two for the Fermi gas in random lattices.
Velický derived the first one within the mean-field, coherent-potential approximation [42]. It holds only for the zero transfer wave vector, $q=0$, and arbitrary frequency
@@
It was derived nonperturbatively, but holds only if the two-particle Hilbert space can be spanned over sums of the products of Bloch waves with real wave vectors $\mathbf{k}$ and $\mathbf{k}^{\prime}$. It means that the two-particle space contains only scattering states and no quantum bound states. This Ward identity guarantees that the wave-function norm is conserved in the Hilbert space of extended states. It prevents a transition to spatially localized states.
@@
Both the Ward identities have a substantial effect on the electron-hole correlation function in the metallic phase. They restrict the low-energy asymptotics to the following form
@@
$n(E)=-\frac{1}{N}\sum_{\mathbf{k}}\int_{-\infty}^{\infty}\frac{d\omega}{\pi}f(\omega)\Im G^{R}_{\mathbf{k}}(E_{F}+\omega)$ (16)
is the density of the electron states at temperature $T$ and energy $E$, and $f(E)$ is the Fermi function. The denominator of the electron-hole correlation function contains the diffusion pole, and its denominator vanishes in the limit $\omega\to 0$ together with $q\to 0$, as demonstrated in the preceding subsection. The linear term in frequency does not depend on the disorder strength due to particle conservation and Ward identity (14). This low-energy expansion is then controlled by two disorder-dependent functions $A(\omega)$ and $D(\omega,\mathbf{q})$. The latter is the generalized diffusion function, and the former time scale was introduced in Ref. [1]. It violates Ward identity Eq. (13) that holds, however, only when no tunneling and bound two-particle states may exist. As argued in Ref. [1], Anderson localization represents a quantum bound state between the moving electron and the hole left behind and $A(\omega)\to\infty$ as $\omega\to 0$, indicating that the tunneling barrier becomes effectively infinite and unpenetrable to infinite distances for the diffusing particle.
The diffusion function $D(\omega,\mathbf{q})$ plays a special role in the description of diffusion. Its homogeneous, static value, diffusion constant $D=D(E;0,\mathbf{0})$, is related to the static electrical conductivity in the Einstein relation, $\sigma=e^{2}nD$. Introducing the inhomogeneous dynamical conductivity $\sigma(E;\omega,\mathbf{q})$, the static Einstein relation can be generalized to a dynamical form using the electron-hole correlation function in the low-energy asymptotics $\omega,q\to 0$:
@@
The disorder-dependent coefficients in the denominator of the electron-hole correlation function in Eq. (15) are proportional to second derivatives $A(E;\omega)\propto\partial^{2}\mathcal{D}(E;\omega,\mathbf{0})/\partial\omega^{2}$ and $D(E;\omega,\mathbf{q})\propto\nabla_{\mathbf{q}}^{2}\mathcal{D}(E;\omega,\mathbf{q})$ as follows from Eq. (12). The eigenvalue $S(E)=1$ and the first frequency derivative $\partial\mathcal{D}(E;\omega,\mathbf{0})/\partial\omega$ evaluated at $\omega=0$ remain disorder-independent as exacted from Ward identity (14).
@@
It is generally assumed that Anderson localization occurs when $D\to 0$ for $n>0$, i.e., when the static diffusion constant vanishes. It is, however, not a critical point. The actual, disorder-induced critical point is when $A\to\infty$. It means that the frequency derivative of the imaginary part of the dynamical conductivity, or the real electric polarizability, diverges. The width of the central peak of the real part of the dynamical conductivity vanishes, but not necessarily its height, the static diffusion. We argue below that this is the critical point of the Anderson localization transition. A new order parameter emerges in the denominator of the electron-hole correlation function beyond this transition. Consequently, the diffusion pole and diffusion vanish as well as the static conductivity $\sigma(E;0)=0$ calculated from the generalized Einstein relation, Eq. (II.3), even if the static diffusion constant from the metallic phase is positive, $D(E)>0$.
It is clear from the exact representation of the electron-hole correlation function, Eqs. (10)-(12), that one cannot derive a closed equation for either parameter $A(E;\omega)$ or $D(E;\omega,\mathbf{q})$. The complete microscopic theory must give, apart from the one-electron self-energy $\Sigma^{R}_{\mathbf{k}}(E)$, expressions for the full irreducible vertex $L^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ and the eigenvector $\phi_{\mathbf{k}}(E)$. The theory of Vollhardt and Wölfle, [33, 34, 37], leading to a self-consistent equation for $D(E;\omega,\mathbf{0})$, is only an effective approximation. Its derivation is not fully microscopically controlled, and it cannot be viewed as a consistent microscopic theory of Anderson localization as long as the corresponding electron-hole irreducible vertex $L^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ remains unknown.
@@ III Perturbation expansion
The complete microscopic theory of charge diffusion and Anderson localization must determine the one-particle and two-particle Green functions in concord with the macroscopic conservation laws. Moreover, it must yield two-particle functions that are not fully determined by the one-particle self-energy for Anderson localization to occur. We hence invert the construction of approximations suitable for describing Anderson localization and employ the bottom-up approach, in which we determine the self-energy from the approximated two-particle irreducible vertex. This will be used beyond the local mean-field theory. The mean-field irreducible vertex, determined from the local self-energy, will serve as input to the two-particle approach.
@@ III.1 Coherent potential approximation
The coherent-potential approximation (CPA) is the best local approximation for the self-energy. It represents a dynamical mean-field theory for the disordered Fermi gas [52]. It is determined for a complex energy $z=E+i\eta$, and $\eta\neq 0$ from a Soven equation [38]
@@
where $\rho(\epsilon)$ is the density of states. It is a single-site approximation with multiple scattering on a single atomic potential $V_{i}$. It does not include back-scattering effects, which arise only when distant scattering centers are included.
The coherent-potential approximation contains all information of the random lattice in the self-energy $\Sigma(z)$. All higher-order irreducible vertices are local and are determined from he self-energy. Consequently, the two-particle irreducible vertex $\lambda(z_{+},z_{-})$ is directly determined from the Ward identity (14)
@@
It is easy to find the eigenvector of the Bethe-Salpeter equation for the electron-hole vertex. It is $\phi_{\mathbf{k}}(E)=G^{R}_{\mathbf{k}}(E)/\sqrt{\left\langle|G^{R}|^{2}\right\rangle}$ and $\left\langle|G^{R}|^{2}\right\rangle=N^{-1}\sum_{\mathbf{k}}\left|G^{R}_{\mathbf{k}}(E)\right|^{2}$ with the eigenvalue $S(E)=1$, since $\lambda(E;0)=1/\left\langle|G^{R}|^{2}\right\rangle$. Representation (12) with this eigenvector gives the exact result of the electron-hole correlation function, Eq. (23).
@@ III.2 Beyond the local mean field
The CPA is a good approximation for the spectral, one-particle properties. It, however, does not go beyond the semi-classical Drude form of the electric conductivity. We must change the strategy in calculating the averaged Green and correlation functions. We replace the self-energy $\Sigma^{R}_{\mathbf{k}}(E)$ by the two-particle irreducible vertex $\Lambda_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ as the fundamental quantity characterizing the selected approximation. The Dyson equation is replaced by a Bethe-Salpeter equation in the electron-hole channel, determining the full two-particle vertex
@@
The electron-hole irreducible dynamical vertex ${\Lambda}^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ will be determined from the renormalized perturbation theory together with the one-particle propagators $G^{RA}_{\mathbf{k}(E)}$. The dynamical vertex ${\Lambda}^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ is different for $\omega\neq 0$ and $q\neq 0$ from the conserving vertex ${L}^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$. Their exact relation will be discussed later. It is important to stress that the dynamical vertex contains more information than the self-energy alůthough it must be related to the one-particle propagators. Only its static homogeneous part, $\omega=0$ and $q=0$ determines the imaginary part of the self-energy as demanded by the static limit of the Ward identity (14)
@@
Equations (25) uniquely determine self-energy $\Sigma^{R}_{\mathbf{k}}(E)$ for the given vertex function $\Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;0,\mathbf{0})$. Unlike the CPA, these equations determine the self-energy only for real energies.
@@ III.3 Parquet equations and two-particle self-consistency
Figure 1: Diagrammatic representation of the Bethe-Salpeter equation for multiple scattering in the electron-hole channel (diffuson). The two-particle vertices, characterized by wave vectors $\mathbf{k},\mathbf{k}^{\prime},\mathbf{q}$, are connected with off-diagonal one-electron propagators, and the double-primed wave vectors are summed over the Brillouin zone.
Figure 2: Diagrammatic representation of the Bethe-Salpeter equation for multiple scattering in the electron-electron channel (cooperon). The two-particle vertices in this channel are characterized by wave vectors $\mathbf{k},\mathbf{k}^{\prime},\mathbf{Q},$. It is connected with the electron-hole channel by inverting one of the electron lines.
An important aspect of diffusion in random media that the CPA misses is backscattering. It can be included only if the approximation contains more than a single scattering centre. The CPA consists of all single-site scattering events and may serve as a suitable starting point for incorporating non-local effects and backscattering. We use the off-diagonal CPA propagator as the unperturbed Green function in the expansion beyond CPA. It is represented in the energy-wave vector space as
@@
The inclusion of non-local scattering yields a crucial improvement over the local CPA. Scattering on spatially distant random potentials enables us to distinguish between the propagation of the particle, diagrammatically represented by an oriented solid line pointing from left to right, and that of the hole, represented by a solid line pointing from right to left. There is no preferred direction for the local CPA propagators. The one-particle functions can be described equivalently in the particle or the hole formalism. The situation is different at the two-particle level. The simultaneous propagation between distinct sites of two particles, shown as parallel lines in the diagrammatic representation, differs from the propagation of the particle-hole pair, shown as antiparallel lines. The two pair propagations define different two-particle irreducibility and generate two non-equivalent representations of the two-particle vertex via Bethe-Salpeter equations with non-local propagators from Eq. (26)[46, 47]. The Bethe-Salpeter integral equations represent multiple scattering in the electron-hole or electron-electron channels and introduce distinct two-particle-irreducible functions. The irreduble vertex $\bar{\Lambda}^{eh}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ in the electron-hole channel differs from the irreducible vertex $\bar{\Lambda}^{ee}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{Q})$ in the electron-electron channel. We denoted the bosonic vector conserved in electron-hole scattering events $\mathbf{q}$ and $\mathbf{Q}=\mathbf{k}+\mathbf{k}^{\prime}-\mathbf{q}$ conserved in the electron-electron channel. The Bethe-Salpeter equations in the electron-hole and electron-electron channels are diagrammatically represented in Figs. 1 and 2, respectively. Their mathematical form is
@@
We distinguished the two-particle vertex according to which Bethe-Salpeter equation determined it, and which conserving bosonic vector we used to characterize it. They are identical, only represented with different transfer vectors. That is, $\Gamma^{eh}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})=\Gamma^{ee}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{k}+\mathbf{k}^{\prime}-\mathbf{q})$.

### v4

@@ II.3 Ward identities, diffusion pole, and generalized Einstein relation
We must know the dynamical electron-hole Green function $\mathcal{G}^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ to determine the electron-hole correlation function and the transport properties of the model. It depends on the one-particle Green functions $G^{R}_{\mathbf{k}}(E_{+})$ and $G^{A}_{\mathbf{k}}(E_{-})$. The solution of the two-particle functions cannot be disconnected from that of the one-particle Green function. The one-particle and two-particle functions are connected via Ward identities. There are two for the Fermi gas in random lattices.
Velický derived the first one within the mean-field, coherent-potential approximation [42]. It holds only for the zero transfer wave vector, $q=0$, and arbitrary frequency.
@@
It was derived nonperturbatively but holds only if the two-particle Hilbert space can be spanned by sums of products of Bloch waves with real wave vectors $\mathbf{k}$ and $\mathbf{k}^{\prime}$. This means that the two-particle space contains only scattering states and no quantum bound states. This Ward identity guarantees that the wave-function norm is conserved in the Hilbert space of extended states. It prevents a transition to spatially localized states.
@@
Both Ward identities have a substantial effect on the electron-hole correlation function in the metallic phase. They restrict the low-energy asymptotics to the following form
@@
$n(E)=-\frac{1}{N}\sum_{\mathbf{k}}\int_{-\infty}^{\infty}\frac{d\omega}{\pi}f(\omega)\frac{\partial}{\partial E}\Im G^{R}_{\mathbf{k}}(E+\omega)$ (16)
is the density of the electron states at temperature $T$ and energy $E$, and $f(E)$ is the Fermi function. The denominator of the electron-hole correlation function contains the diffusion pole, and it vanishes in the limit $\omega\to 0$ together with $q\to 0$, as demonstrated in the preceding subsection. The linear term in frequency does not depend on the disorder strength due to particle conservation and the Ward identity (14). This low-energy expansion is then controlled by two disorder-dependent functions $A(\omega)$ and $D(\omega,\mathbf{q})$. The latter is the generalized diffusion function, and the former time scale was introduced in Ref. [1]. It violates the Ward identity Eq. (13), which holds, however, only when no tunneling and bound two-particle states may exist. As argued in Ref. [1], Anderson localization represents a quantum bound state between the moving electron and the hole left behind, and $A(\omega)\to\infty$ as $\omega\to 0$, indicating that the tunneling barrier becomes effectively infinite and unpenetrable to infinite distances for the diffusing particle.
The diffusion function $D(\omega,\mathbf{q})$ plays a special role in the description of diffusion. Its homogeneous, static value, the diffusion constant $D=D(E;0,\mathbf{0})$, is related to the static electrical conductivity through the Einstein relation, $\sigma=e^{2}nD$. Introducing the inhomogeneous dynamical conductivity $\sigma(E;\omega,\mathbf{q})$, the static Einstein relation can be generalized to a dynamical form using the electron-hole correlation function in the low-energy asymptotics $\omega,q\to 0$:
@@
The disorder-dependent coefficients in the denominator of the electron-hole correlation function in Eq. (15) are proportional to second derivatives $A(E;\omega)\propto\partial^{2}\mathcal{D}(E;\omega,\mathbf{0})/\partial\omega^{2}$ and $D(E;\omega,\mathbf{q})\propto\nabla_{\mathbf{q}}^{2}\mathcal{D}(E;\omega,\mathbf{q})$, as follows from Eq. (12). The eigenvalue $S(E)=1$ and the first frequency derivative $\partial\mathcal{D}(E;\omega,\mathbf{0})/\partial\omega$ evaluated at $\omega=0$ remain disorder-independent, as exacted from the Ward identity (14).
@@
Anderson localization is generally assumed to occur when $D\to 0$ for $n>0$, i.e., when the static diffusion constant vanishes. However, this is not a critical point. The disorder-induced critical point occurs when $A\to\infty$. This means that the frequency derivative of the imaginary part of the dynamical conductivity, or the real electric polarizability, diverges. The width of the central peak of the real part of the dynamical conductivity vanishes, but not necessarily its height, the static diffusion. We argue below that this is the critical point of the Anderson localization transition. Beyond this transition, a new order parameter emerges in the denominator of the electron-hole correlation function. Consequently, the diffusion pole and diffusion vanish, as does the static conductivity $\sigma(E;0)=0$ calculated from the generalized Einstein relation, Eq. (II.3), even if the static diffusion constant from the metallic phase is positive, $D(E)>0$.
It is clear from the exact representation of the electron-hole correlation function, Eqs. (10)-(12), that one cannot derive a closed equation for either parameter $A(E;\omega)$ or $D(E;\omega,\mathbf{q})$. The complete microscopic theory must give, apart from the one-electron self-energy $\Sigma^{R}_{\mathbf{k}}(E)$, expressions for the full irreducible vertex $L^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ and the eigenvector $\phi_{\mathbf{k}}(E)$. The theory of Vollhardt and Wölfle, [33, 34, 37], which leads to a self-consistent equation for $D(E;\omega,\mathbf{0})$, is only an effective approximation. Its derivation is not fully microscopically controlled, and it cannot be viewed as a consistent microscopic theory of Anderson localization as long as the corresponding electron-hole irreducible vertex $L^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ remains unknown.
@@ III Perturbation expansion
The complete microscopic theory of charge diffusion and Anderson localization must determine the one-particle and two-particle Green functions in accordance with the macroscopic conservation laws. Moreover, it must yield two-particle functions that are not fully determined by the one-particle self-energy for Anderson localization to occur. We hence invert the construction of approximations suitable for describing Anderson localization and employ a bottom-up approach, in which we determine the self-energy from the approximated two-particle irreducible vertex. This will be used beyond the local mean-field theory. The mean-field irreducible vertex, determined from the local self-energy, will serve as input to the two-particle approach.
@@ III.1 Coherent potential approximation
The coherent-potential approximation (CPA) is the best local approximation to the self-energy. It represents a dynamical mean-field theory for the disordered Fermi gas [52]. It is determined for a complex energy $z=E+i\eta$ with $\eta\neq 0$ from a Soven equation [38].
@@
where $\rho(\epsilon)$ is the density of states. It is a single-site approximation with multiple scattering on a single atomic potential $V_{i}$. It does not include backscattering effects, which arise only when distant scattering centers are included.
The coherent-potential approximation contains all information about the random lattice in the self-energy $\Sigma(z)$. All higher-order irreducible vertices are local and are determined from the self-energy. Consequently, the two-particle irreducible vertex $\lambda(z_{+},z_{-})$ is determined directly from the Ward identity (14).
@@
It is easy to find the eigenvector of the Bethe-Salpeter equation for the electron-hole vertex. It is $\phi_{\mathbf{k}}(E)=G^{R}_{\mathbf{k}}(E)/\sqrt{\left\langle|G^{R}|^{2}\right\rangle}$, with eigenvalue $S(E)=1$, since $\lambda(E;0)=1/\left\langle|G^{R}|^{2}\right\rangle$. The eigenvalue is $S(E)=1$, and the eigenvector is $\phi_{\mathbf{k}}(E)=G^{R}_{\mathbf{k}}(E)/\sqrt{\left\langle|G^{R}|^{2}\right\rangle}$. The representation (12) with this eigenvector gives the exact result for the electron-hole correlation function, Eq. (23).
@@ III.2 Beyond the local mean field
The CPA is a good approximation for spectral one-particle properties. It, however, does not go beyond the semiclassical Drude form of the electric conductivity. We must change the strategy for calculating the averaged Green and correlation functions. We replace the self-energy $\Sigma^{R}_{\mathbf{k}}(E)$ by the two-particle irreducible vertex $\Lambda_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ as the fundamental quantity characterizing the selected approximation. The Dyson equation is replaced by a Bethe-Salpeter equation in the electron-hole channel, which determines the full two-particle vertex.
@@
The electron-hole irreducible dynamical vertex ${\Lambda}^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ will be determined from renormalized perturbation theory together with the one-particle propagators $G^{RA}_{\mathbf{k}(E)}$. The dynamical vertex ${\Lambda}^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ differs from the conserving vertex ${L}^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ when $\omega\neq 0$ or $q\neq 0$. Their exact relation will be discussed later. It is important to stress that the dynamical vertex contains more information than the self-energy, although it must be related to the one-particle propagators. Only its static homogeneous part, $\omega=0$ and $q=0$, determines the imaginary part of the self-energy, as demanded by the static limit of the Ward identity (14)
@@
Equations (25) uniquely determine the self-energy $\Sigma^{R}_{\mathbf{k}}(E)$ for the given vertex function $\Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;0,\mathbf{0})$. Unlike the CPA, these equations determine the self-energy only for real energies.
@@ III.3 Parquet equations and two-particle self-consistency
Figure 1: Diagrammatic representation of the Bethe-Salpeter equation for multiple scattering in the electron-hole channel (diffuson). The two-particle vertices, labeled by wave vectors $\mathbf{k},\mathbf{k}^{\prime},\mathbf{q}$, are connected by off-diagonal one-electron propagators, and the double-primed wave vectors are summed over the Brillouin zone.
Figure 2: Diagrammatic representation of the Bethe-Salpeter equation for multiple scattering in the electron-electron channel (cooperon). The two-particle vertices in this channel are characterized by wave vectors $\mathbf{k},\mathbf{k}^{\prime},\mathbf{Q}$. It is connected to the electron-hole channel by inverting one of the electron lines.
An important aspect of diffusion in random media that the CPA misses is backscattering. It can be included only if the approximation contains more than a single scattering center. The CPA accounts for all single-site scattering events and may serve as a suitable starting point for incorporating non-local effects and backscattering. We use the off-diagonal CPA propagator as the unperturbed Green function in the expansion beyond CPA. It is represented in the energy-wave vector space as
@@
Inclusion of non-local scattering yields a crucial improvement over the local CPA. Scattering on spatially distant random potentials enables us to distinguish between the propagation of the particle, diagrammatically represented by an oriented solid line pointing from left to right, and that of the hole, represented by a solid line pointing from right to left. There is no preferred direction for the local CPA propagators. The one-particle functions can be described equivalently in the particle or the hole formalism. The situation is different at the two-particle level. The simultaneous propagation of two particles between distinct sites, shown as parallel lines in the diagrammatic representation, differs from the propagation of the particle-hole pair, shown as antiparallel lines. The two pair propagations define different two-particle irreducibility and generate two non-equivalent representations of the two-particle vertex via Bethe-Salpeter equations with non-local propagators from Eq. (26)[46, 47]. The Bethe-Salpeter integral equations represent multiple scattering in the electron-hole or electron-electron channels and introduce distinct two-particle-irreducible functions. The irreducible vertex $\bar{\Lambda}^{eh}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ in the electron-hole channel differs from the irreducible vertex $\bar{\Lambda}^{ee}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{Q})$ in the electron-electron channel. We denote the bosonic vector conserved in electron-hole scattering events $\mathbf{q}$ and in the electron-electron channel $\mathbf{Q}=\mathbf{k}+\mathbf{k}^{\prime}-\mathbf{q}$. The Bethe-Salpeter equations in the electron-hole and electron-electron channels are diagrammatically represented in Figs. 1 and 2, respectively. Their mathematical form is
@@
We distinguished the two-particle vertices by the Bethe-Salpeter equation that determined them and by the conserving bosonic vectors we used to characterize them. The two equations are identical, differing only in the transfer vectors. That is, $\Gamma^{eh}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})=\Gamma^{ee}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{k}+\mathbf{k}^{\prime}-\mathbf{q})$.